I need help calculating c, cu, cm, and alpha for a vane less diffuser assuming no diffuser losses. My numbers for the outlet of the turbo are U=23.561 m/s
Cm=2.8318 m/s
Cu= 1.237 m/s
r=. 45m

Cu is absolute velocity that is tangential
Cm is effective volume flow
C is magnitude of those
Alpha is the angle of the flow
Answer #1

Answer:

The solution is given below in the fig.

Given:- U = 23.561 m/s Cm= 2.8318 m/s C4 = 1.237 m/s. = 0.4 sm 63 Cm Q-flow angle is calculated by tana = Cm - 2.8318 I 1.237